An air bag increase impact time to reduce force experienced by a driver of a car during collision. A bus hits a car that is parked on the road from behind. The car accelerates to 4.5 m/s in 0.15 s . what force does the driver of the car experience if he is not wearing the seatbelt during the accident . Assume that the combine mass of the car and the driver is 2580 kg . if the car has an air bag that increased the impact time to 0.450 s, what force was experience by the driver ?
1
Expert's answer
2020-12-06T17:29:48-0500

First, write the equation for acceleration:


"a=\frac vt."


Find force by Newton's second law:


"F=ma=m\frac vt."


Without the airbag, the force will be


"F_1=2850\frac {4.5}{0.15}=85500\text{ N}.\\\space\\
F_2=2850\frac {4.5}{0.45}=28500\text{ N}"

with the airbag. We see that the airbag reduces the force thrice.